OpenStudy (jkristia):

Hi, I have a problem understanding the solution to the problem in 1C-2. In the solution, how come f(a) becomes 0 ? 1C-2 Let f(z) = (z -a)g(z). Use the definition of the derivative to calculate that f'(a) = g(a).

OpenStudy (jkristia):

I got the answer from another forum. f(a) becomes 0 because x -> a, meaning f(a) can be written as (x-a)g(a) = (a - a)g(a) = 0*g(a) = 0.